OpenStudy (anonymous):

A man in a train traveling at 30 km/hr notices that a train going in the opposite direction passes him in 9 seconds. If the length of this train is 200m, find its speed?

OpenStudy (anonymous):

solve \[200=(30+r)9\] for r

OpenStudy (anonymous):

distance=rate x time The train's relative speed is 200/9 m/s=80 km/hr Speed=80-30=50 km/hr

OpenStudy (siddharth):

Let the other train's speed be v km/hr. Now, if we consider the man to be stationary (his frame of reference), the other train will be travelling at 30+v km/hr in the opposite direction. So, other train travels 200m in 9 seconds. Distance = speed * time 200/1000 km = (30+v) km/hr * 9/3600 hrs 20 = (30+v)/4 v = 50 km /hr

OpenStudy (anonymous):

Hello, siddharth, but 200/1000 is 0.2

OpenStudy (siddharth):

I did some simplification: \[\frac{200}{1000} = \frac{9(30+v)}{3600}\] multiply both sides by 1000: \[200 = \frac{9(30+v)}{3600} \times 1000\] \[200 = \frac{9(30+v)}{36} \times 10\] divide both sides by 10 \[20 = \frac{9(30+v)}{36} \]